# Predictive Analysis for Optimizing Port Operations

### v1

@@ Abstract
Maritime transport is a pivotal logistics mode for the long-distance and bulk transportation of goods. However, the intricate planning involved in this mode is often hindered by uncertainties, including weather conditions, cargo diversity, and port dynamics, leading to increased costs. Consequently, accurately estimating vessel total (stay) time at port and potential delays becomes imperative for effective planning and scheduling in port operations. This study aims to develop a port operation solution with competitive prediction and classification capabilities for estimating vessel Total and Delay times. This research addresses a significant gap in port analysis models for vessel Stay and Delay times, offering a valuable contribution to the field of maritime logistics. The proposed solution is designed to assist decision-making in port environments and predict service delays. This is demonstrated through a case study on Brazil ports. Additionally, feature analysis is used to understand the key factors impacting maritime logistics, enhancing the overall understanding of the complexities involved in port operations.
@@ 1 Introduction
In the intricate landscape of global maritime logistics, an often overlooked yet substantial segment of a vessel’s journey unfolds within the confines of port facilities. While traditional narratives tend to focus on vessels (ships) navigating the open waters [Ogura et al., 2021, Kolley et al., 2022, Mekkaoui et al., 2022], a substantial portion of their journey involves intricate maneuvers, cargo handling, and operational intricacies within ports [Nguyen et al., 2019]. Surprisingly, this dualistic nature of a vessel’s journey—both at sea and within port—has received scant attention in the existing literature. The temporal dynamics, efficiency challenges, and optimization opportunities inherent in the port-centric leg of a vessel’s journey warrant closer scrutiny [Abreu et al., 2022b]. This gap in scholarly exploration calls for an in-depth examination, shedding light on the pivotal role that port activities play in the broader context of maritime transportation, supply chain efficiency, and global trade networks [Dimitrios et al., 2023].
The optimization of port operations holds critical importance in the context of globalization, where nations strive to maintain their competitiveness and economic resilience in an interconnected world. Efficient port operations are essential not only for facilitating trade and commerce but also for accommodating the growing flow of goods and passengers associated with global tourism and cultural exchanges. The integration of marine traffic information through APIs has emerged as a key technological advancement, offering real-time insights into vessel movements and enabling proactive decision-making to enhance port efficiency on a global scale.
While the specific focus of this project is on Brazil ports, the broader implications extend to a global context. Ports worldwide face similar challenges related to the optimization of operations, the impact of globalization on trade flows, and the integration of advanced technologies for enhanced efficiency [Çağatay Iris and Lam, 2019, Izaguirre et al., 2020a]. By studying the case of Brazil, with its rich maritime history and strategic positioning in global trade, valuable insights can be gained that are applicable to port operations and maritime logistics across diverse regions and economies.
@@
  • We define the problem for port operational analysis which is lacking in the current literature.
  • We execute an exhaustive comparison of methods with multiple metrics for prediction and classification of Total time (stay time) and Delay time at port.
@@
The paper is structured as follows: Section 2 provides a comprehensive review of the literature relevant to the study. Section 3 discusses the Brazil port case study and outlines the proposed methods used for port analysis. Section 4 dives into the results and share keys features that impact the port operations. We finish with Section 5 that offers the conclusions from the study, and suggests avenues for future research.
@@ 2 Literature Review
The Berth Allocation Problem (BAP) is a critical challenge in port operations, aiming to optimize the allocation of berthing facilities for the incoming ships while considering various restrictions and objectives [Bierwirth and Meisel, 2015]. The estimation of different durations for the vessels is a crucial factor in BAP mathematical formulations, as it directly influences berth allocation decisions. Therefore, the development of efficient prediction and classification models for Total time is strategically important for addressing the BAP. Studies [Guo et al., 2023, Rodrigues and Agra, 2022] have explored novel algorithms and optimization techniques to address the BAP, emphasizing the importance of accurate Total time predictions in optimizing berth allocations.
Predictive Model Scheduling (MPS) is another area of interest in port operations, focusing on the development of predictive models to optimize scheduling processes for port activities [Dulebenets, 2018, van Boetzelaer, 2013]. By leveraging predictive analytics, MPS aims to enhance the efficiency and effectiveness of port operations by forecasting demand, optimizing resource allocation, and improving overall performance [Weerasinghe et al., 2023]. These works have explored the application of machine learning algorithms for predictive scheduling in container terminals, demonstrating the potential of these methods in improving port productivity.
@@
Despite the significant work on port operations, there is a absence of studies on models for predicting and classifying Total time and Delay time of vessels at ports, which highlights the need for research in this area. By addressing this gap, the proposed study aims to contribute valuable insights that can inform decision-making processes in port management, ultimately improving the efficiency and effectiveness of port operations. This information will be crucial for solving other port related problems such as the BAP, as accurate predictions of Total time and Delay time can enhance the applicability of BAP solutions in real-world port environments. Moreover, these insights can support management decisions related to the planning, allocation, and scheduling of loading and unloading operations, contributing to overall operational excellence in port logistics.
@@ 3 Case Study and Methodology
The Case study of Brazil port is from [Abreu et al., 2022b, Abreu et al., 2022a]. Brazil’s marine logistics and port operations are crucial components of its economy due to the country’s extensive coastline and reliance on maritime transport for international trade. Brazil boasts one of the world’s most extensive coastlines, stretching over 7,400 kilometers, with numerous ports facilitating the import and export of goods. The country’s ports play a pivotal role in connecting Brazil to global markets, serving as crucial gateways for international trade [Cabral and Ramos, 2018, López-Bermúdez et al., 2019]. However, port logistics in Brazil are complex due to inflexible regulations and the need for extensive information, impacting delivery times and costs [da Veiga Lima and de Souza, 2022, Galvao et al., 2017]. Maritime logistics and port operations are vital for Brazil’s economy for several reasons: Trade Facilitation, Economic Growth, Job Creation, Transportation Network, and Strategic Importance.
Overall, the efficient functioning of Brazil’s marine logistics and port operations is essential for the country’s economic prosperity, trade competitiveness, and integration into the global economy [Costa et al., 2022, Cacho et al., 2021]. It represents a critical component of Brazil’s infrastructure and contributes significantly to its overall development.
@@
Finally, the stay (Total) time characteristics section includes variables related to the duration of a vessel’s stay in port. This encompasses details such as the arrival date and hour of the ship, the mooring date and hour, the unblocking date and hour, the start and end dates and hours of operations, the month of mooring, and various time durations (T1-Wait time for the moat, T2-Wait time for start of operation, T3-Operation time, T4-Wait time for unberthing, T2+T3+T4-Mooring time, and T1+T2+T3+T4-Total time) that reflect different aspects of the vessel’s stay in port. These variables are critical for understanding the temporal aspects of port operations, including the time spent waiting, in operation, and overall stay duration. We define Delay time as Total time-Operation time (T1+T2+T4).
We define the Total time (stay time) class using Fig. 1, similar to [Abreu et al., 2022b]. Class 0 (Low) is defined for vessels with Total time at port for 0-25 hours, Class 1 (Medium) is defined for 25-50 hours, Class 2 (High) includes 50-75 hours and Class 3 (Very High) is $>$75 hours. For Delay Class, we use the plots as well as summary statistics (mean Delay $\approx$ 10 hours, variance of Delay $\approx$125), and define class 0 as Delay time $<$ 24 hours and Class 1 as Delay time $>$ 24 hours.
@@
In the realm of supervised learning for prediction and classification, tree-based methods and statistics-based methods represent two prominent approaches, each with its strengths and applications. Tree-based algorithms, such as Random Forest, Extreme Gradient Boosting (XGBoost), AdaBoost, Light Gradient Boosting Machine (LightGBM), and Extra Trees, are known for their ability to handle complex relationships in data and provide robust predictions. Random Forest, for instance, constructs multiple decision trees and aggregates their outputs to improve accuracy and reduce overfitting [Breiman, 2001]. XGBoost, a gradient boosting algorithm, excels in optimizing complex objective functions and has been widely used in competitions and industry applications for its performance [Chen and Guestrin, 2016]. AdaBoost focuses on iteratively boosting the performance of weak learners to achieve strong predictive models [Freund and Schapire, 1997]. LightGBM is known for its efficiency in handling large datasets and has become popular for its speed and accuracy [Ke et al., 2017]. Extra Trees, or Extremely Randomized Trees, is an ensemble learning method that builds multiple decision trees and averages their predictions, providing robustness against noise and overfitting [Geurts et al., 2006]. Additionally, feature importance analysis in tree-based models helps identify the most influential variables in the prediction process, providing insights into the underlying patterns in the data [Louppe et al., 2013].
@@
On the other hand, statistics-based methods like regression, Support Vector Machines (SVM), Bayesian classifiers, k-Nearest Neighbors (kNN), and Elastic Net offer a different set of tools for prediction and classification tasks. Regression techniques, including linear regression and its variants, are widely used for predicting continuous outcomes based on the relationships between variables [James et al., 2013]. SVM, a powerful classification method, aims to find the optimal hyperplane that best separates data points into different classes in a high-dimensional space [Cortes and Vapnik, 1995]. Bayesian classifiers, based on Bayes’ theorem, are probabilistic models that calculate the probability of a sample belonging to a particular class [Hand and Yu, 2001]. kNN, a simple yet effective algorithm, classifies a data point based on the majority class of its nearest neighbors in the feature space [Altman, 1992]. Elastic Net, a regularization technique combining L1 (Lasso) and L2 (Ridge) penalties, is particularly useful for feature selection and dealing with multicollinearity in regression problems [Zou and Hastie, 2005]. Each of these statistics-based methods has its own assumptions, strengths, and limitations, making them suitable for different types of datasets and prediction tasks.
@@
In both prediction and classification tasks, various metrics are used to evaluate the performance of machine learning models. For prediction tasks, common metrics include Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), Mean Squared Error (MSE), R-squared ($R^{2}$), Root Mean Squared Logarithmic Error (RMSLE), and Mean Absolute Percentage Error (MAPE). MAE measure the average magnitude of the errors between predicted and actual values. MSE is the average of the squared differences between predicted and actual values (RMSE is square root of MSE). $R^{2}$ represents the proportion of the variance in the dependent variable that is predictable from the independent variables. RMSLE is particularly useful for tasks where the target variable spans several orders of magnitude. MAPE measures the percentage difference between predicted and actual values. For all the metrics, lower values indicate better fit whereas for $R^{2}$ values close to 1 are considered good.
@@
For classification tasks, metrics such as Accuracy, Precision, Recall, Area Under the Receiver Operating Characteristic Curve (AUC), F1 score, and Cohen’s Kappa are commonly used. All this values can be derived from the confusion matrix. Accuracy measures the proportion of correctly classified instances over the total number of instances. Precision measures the proportion of true positive predictions among all positive predictions, focusing on the accuracy of positive predictions. Recall (also known as sensitivity) measures the proportion of true positive predictions among all actual positives, focusing on the model’s ability to detect positive instances. AUC measures the model’s ability to distinguish between classes. F1 score is the harmonic mean of precision and recall, providing a balance between the two metrics. Cohen’s Kappa measures the agreement between predicted and actual classifications, accounting for the possibility of the agreement occurring by chance. For all the metrics, higher values indicate good classification performance.
@@
In the context of port operations, the aforementioned machine learning methods are applied to predict and classify Total time and Delay time, which are critical factors in port logistics. By leveraging these techniques, we gain valuable insights into the complex dynamics of port operations.
@@ 4 Results
We split the data (n=153331) into 80% (122664) for training and 20% (30667) for testing. We consider only 24 features and remove any feature information related to calculation of Total time and Delay time at the port. To ensure we don’t over fit and to pick the best hyperparameters, we perform 10-fold cross validation (CV). We first look into the prediction of Total time and Delay time, followed by classification of Total time and Delay time. The experiments were performed on an 8-core M2 CPU.
Also, we perform feature analysis for a comprehensive understanding of the key factors influencing port operations. Addressing these factors is crucial for optimizing the Berth Allocation Problem (BAP), Predictive Model Scheduling (MPS), port performance management, and risk assessment. By enhancing these aspects, ports can achieve greater sustainability, aligning with the goal of promoting green practices.
Table 1 shows the performance of different Machine Learning approaches for Total time prediction. We see a general trend that tree based methods are performing better than statistical approaches on this task. This might be because of the non linear nature in the data. We observe that Random Forest (RF) performs the best, has the lowest RMSE (10.4377) and also has superior values in other metrics in comparison to rest of the methods. Performance on the test set for RF is similar, where the RMSE is 10.2577 (other metrics are also in close range). This is visible from Fig. 3, where the predicted and true values are near to the identity line with $R^{2}$ value of 0.8287.
Fig. 2 shows the important features for the Total time prediction, and we can observe that berth feature plays the most crucial role due to some berths being more busy and less efficient compared to others, followed by other features like moved value, VL Containerized Load Weight, type of mooring navigation, source, destination, month, etc. Berths play a vital role in port operations and hence, efficient berth management is essential for optimizing port capacity, throughput, and overall operational efficiency.
@@
Figure 2: Feature importance for of Total time prediction at port
Figure 3: R-squared plot on test data for Total time prediction at port
We see a similar story from Table 2 where the tree based methods are performing better for predicting the Delay time at port. Random Forest (RF) again performs the best, with an RMSE of 4.2179. The test set performance for RF gives a RMSE of 4.3707 (other metrics are also in similar range). From Fig. 3, we can see that the predicted and true values are not close to the identity line ($R^{2}$ is 0.6086). The prediction of Delay time might be limited by the feature set, as it is known that weather information, and other external factors play a key role with respect to delay.
Fig. 4 shows the important features for the Delay time prediction, and we can observe that moved value, VL Containerized Load Weight, and Berth are key features followed by other features like source, destination, month, etc. It is intuitive that vessels with higher moved value, and VL Containerized Load Weight goes through longer operations leading to build up in Delay time.
@@
Figure 4: Feature importance for of Delay time prediction at port
Figure 5: R-squared plot on test data for Delay time prediction at port
In the classification task, the performance of various Machine Learning methods for (Multi-Class) classification of Total time is depicted in Table 3. A consistent trend is observed where tree-based methods outperform statistical approaches, potentially due to the complex nature of the data. Specifically, Extreme Gradient Boosting (XGBoost) demonstrates the best performance, achieving the highest Accuracy (0.8424) and superior values across other metrics compared to alternative methods. This is 10% higher than the result in [Abreu et al., 2022b]. The test set results further confirm XGBoost’s efficacy, with an Accuracy of 0.8415 and other metrics exhibiting similarly favorable performance as derived from the confusion matrix in Fig. 7. Also, we observe that is it easier to detect the outer classes, Class 0 (Low) and Class 3 (Very High), compared to the other two.
Additionally, Fig. 6 showcases the key features influencing Total time classifcation. Notably, the geographical region (north), Port Complex (Manaus), Type of Mooring Navigation (inside) emerges as the most influential features, followed by factors such as City, Mooring port, Inland transportation route, type of mooring navigation (Long Haul), Berth, etc. The significance of the geographical features is related to the huge coastline of Brazil and the complex dynamics of port operations.
@@
Figure 6: Feature importance for of Total time classification at port
Figure 7: Confusion matrix on test data for Total time classification at port
Table 4 presents the binary classification performance on Delay time. The tree-based methods are again outperforming statistical approaches. XGBoost gives the best performance by achieving the highest Accuracy (0.9511) and best values across other metrics compared to other methods. XGBoost achieves an Accuracy of 0.9505 on the test set and other metrics showing close performance, as illustrated by the confusion matrix in Fig. 9.
Furthermore, Fig. 8 highlights the key features influencing Delay time classification, factors such as Flag Long Course, Flag Cabotage, Owner’s Nationality and type of mooring navigation (Long Haul) emerge as the most influential. This is followed by variables like Port Complex, Load Operation Type, ST Cargo Nature, and Month. Features related to operation characteristics are having the most impact on Delay class.

### v2

@@ Abstract
Maritime transport is a pivotal logistics mode for the long-distance and bulk transportation of goods. However, the intricate planning involved in this mode is often hindered by uncertainties, including weather conditions, cargo diversity, and port dynamics, leading to increased costs. Consequently, accurate estimation of the total (stay) time of the vessel and any delays at the port are essential for efficient planning and scheduling of port operations. This study aims to develop predictive analytics to address the shortcomings in the previous works of port operations for a vessel’s Stay Time and Delay Time, offering a valuable contribution to the field of maritime logistics. The proposed solution is designed to assist decision-making in port environments and predict service delays. This is demonstrated through a case study on Brazil’s ports. Additionally, feature analysis is used to understand the key factors impacting maritime logistics, enhancing the overall understanding of the complexities involved in port operations. Furthermore, we perform SHapley Additive exPlanations (SHAP) analysis to interpret the effects of the features on the outcomes and understand their impact on each sample, providing deeper insights into the factors influencing port operations.
@@ 1 Introduction
In the intricate landscape of global maritime logistics, an often overlooked yet substantial segment of a vessel’s journey unfolds within the confines of port facilities. While traditional narratives tend to focus on vessels (ships) navigating the open waters [Ogura et al., 2021, Kolley et al., 2022a, Mekkaoui et al., 2022], a substantial portion of their journey involves complex maneuvers, cargo handling, and operational intricacies within ports [Nguyen et al., 2019]. Surprisingly, this dualistic nature of a vessel’s journey, at sea, and within port, the latter has received scant attention in the existing literature. The temporal dynamics, efficiency challenges, and optimization opportunities inherent in the port-centric leg of a vessel’s journey warrant closer scrutiny [Abreu et al., 2022]. This gap in scholarly exploration calls for an in-depth examination, shedding light on the pivotal role that port activities play in the broader context of maritime transportation, supply chain efficiency, and global trade networks [Dimitrios et al., 2023].
The optimization of port operations holds critical importance in the context of globalization, where nations strive to maintain their competitiveness and economic resilience in an interconnected world. Efficient port operations are essential not only for facilitating trade and commerce but also for accommodating the growing flow of goods and passengers associated with global tourism and cultural exchanges. The integration of marine traffic information through Application Programming Interface (API) [Reggiannini et al., 2019, Bautista-Sánchez et al., 2019, Eriksen et al., 2006] has emerged as a key technological advancement, offering real-time insights into vessel movements and enabling proactive decision-making to enhance port efficiency on a global scale.
While the specific focus of this project is on Brazil’s ports, the broader implications extend to a global context. Ports worldwide face similar challenges related to the optimization of operations, the impact of globalization on trade flows, and the integration of advanced technologies for enhanced efficiency [Çağatay Iris and Lam, 2019, Izaguirre et al., 2020a]. By studying the case of Brazil, with its rich maritime history and strategic positioning in global trade, valuable insights can be gained that are applicable to port operations and maritime logistics across diverse regions and economies.
@@
  • We define the problem for predictive analysis for port operations which is lacking in the current literature.
  • We execute an exhaustive comparison of methods with multiple metrics for prediction and classification of Total Time (stay time) and Delay Time at the port.
@@
  • We perform SHAP (SHapley Additive exPlanations) analysis to understand the contribution of these key factors on the output.
The paper is structured as follows: Section 2 provides a comprehensive review of the literature relevant to the study. Section 3 discusses Brazil’s port case study and outlines the proposed methods used for port predictive analysis. Section 4 dives into the results and shares key features which impact the port operations. We finish with Section 5, which offers the conclusions from the study, and suggests avenues for future research.
@@ 2 Literature Review
The Berth Allocation Problem (BAP) is a critical challenge in port operations, aiming to optimize the allocation of berthing facilities for the incoming ships while considering various restrictions and objectives [Bierwirth and Meisel, 2015, Kolley et al., 2022b]. The estimation of different durations for the vessels is a crucial factor in BAP mathematical formulations, as it directly influences berth allocation decisions. Therefore, the development of efficient prediction and classification models for Total Time is strategically important for addressing the BAP. Studies [Guo et al., 2023, Rodrigues and Agra, 2022] have explored novel algorithms and optimization techniques to address the BAP, emphasizing the importance of accurate Total Time predictions in optimizing berth allocations.
Predictive Model Scheduling (PMS) is another area of interest in port operations, focusing on the development of predictive models to optimize scheduling processes for port activities [Dulebenets, 2018, van Boetzelaer et al., 2014]. By leveraging predictive analytics for Total Time and Delay Time, PMS aims to enhance the efficiency and effectiveness of port operations by forecasting demand, optimizing resource allocation, and improving overall performance [Weerasinghe et al., 2023]. These works have explored the application of machine learning algorithms for predictive scheduling in container terminals, demonstrating the potential of these methods in improving port productivity.
@@
Despite the significant work on port operations, there is an absence of studies on models for predicting and classifying Total Time and Delay Time of vessels at ports, which highlights the need for research in this area [Chu et al., 2024, Štepec et al., 2020]. By addressing this gap, the proposed study aims to contribute insightful information for the decision-making processes in port management, ultimately improving the efficiency and effectiveness of port operations. This information will be crucial for solving other port related problems such as the BAP, as accurate predictions of Total Time and Delay Time can enhance the applicability of BAP solutions in real-world port environments. Moreover, these insights can support management decisions related to the planning, allocation, and scheduling of loading and unloading operations, contributing to overall operational excellence in port logistics.
Prediction and classification both play important roles in this context. Classification is a relatively easier task to complete and is useful for implementing measures that impact operations on a global scale. In contrast, prediction, though more challenging, is more valuable and precise, as it can help improve operations by addressing minute details. A similar analogy applies to Total Time and Delay Time, as they are approximate values and depend on different operations and the information available. Depending on the modeling task and the variables used, different features impact them to varying extents, providing valuable insights in the predictive analysis of port operations. The combination of these approaches ensures a comprehensive strategy, where classification enables broad decision-making and prediction drives detailed, data-driven improvements. By addressing these dimensions, the study offers a well-rounded approach to enhancing port operations.
@@ 3 Case Study and Methodology
The Case study of Brazil’s ports is from [Abreu et al., 2022]. Brazil’s marine logistics and port operations are crucial components of its economy due to the country’s extensive coastline and reliance on maritime transport for international trade. Brazil boasts one of the world’s most extensive coastlines, stretching over 7,400 kilometers, with numerous ports facilitating the import and export of goods. The country’s ports play a pivotal role in connecting Brazil to global markets, serving as crucial gateways for international trade [Cabral and Ramos, 2018, López-Bermúdez et al., 2019]. However, port logistics in Brazil are complex due to inflexible regulations and the need for extensive information, impacting delivery times and costs [da Veiga Lima and de Souza, 2022, Galvao et al., 2017]. Overall, the efficient functioning of Brazil’s marine logistics and port operations is essential for the country’s economic prosperity, trade competitiveness, and integration into the global economy [Costa et al., 2022, Cacho et al., 2021]. It represents a critical component of Brazil’s infrastructure and contributes significantly to its overall development.
@@
Finally, the Stay (Total) time characteristics section includes variables related to the duration of a vessel’s stay in port. This encompasses details such as the arrival date and hour of the ship, the mooring date and hour, the unblocking date and hour, the start date, end date and hours of operations, the month of mooring, and various time durations (T1-Wait time for the moat, T2-Wait time for start of operation, T3-Operation time, T4-Wait time for unberthing, T2+T3+T4-Mooring time, and T1+T2+T3+T4-Total Time) that reflect different aspects of the vessel’s stay at port. These variables are critical for understanding the temporal aspects of port operations, including the time spent waiting, in operation, and overall stay duration. We define Delay Time as Total Time-Operation time (T1+T2+T4).
[…]
@@
In the realm of supervised learning for prediction and classification, tree-based methods, statistics-based methods, and deep learning represent three prominent approaches, each with its own strengths and applications. Tree-based algorithms, such as Random Forest, Extreme Gradient Boosting (XGBoost), AdaBoost, Light Gradient Boosting Machine (LightGBM), and Extra Trees, are known for their ability to handle complex relationships in data and provide robust predictions. Random Forest, for instance, constructs multiple decision trees and aggregates their outputs to improve accuracy and reduce overfitting [Breiman, 2001]. XGBoost is a gradient boosting algorithm that excels in optimizing complex objective functions and has been widely used in competitions and industry applications for its performance [Chen and Guestrin, 2016]. AdaBoost focuses on iteratively boosting the performance of weak learners to achieve strong predictive models [Freund and Schapire, 1997]. LightGBM is known for its efficiency in handling large datasets and has become popular for its speed and accuracy [Ke et al., 2017]. Extra Trees or Extremely Randomized Trees is an ensemble learning method that builds multiple decision trees and averages their predictions, providing robustness against noise and overfitting [Geurts et al., 2006]. Additionally, feature importance analysis in tree-based models helps identify the most influential variables in the prediction process, providing insights into the underlying patterns in the data [Louppe et al., 2013].
@@
On the other hand, statistics-based methods like regression, Support Vector Machines (SVM), Bayesian classifiers, k-Nearest Neighbors (kNN), and Elastic Net offer a different set of tools for prediction and classification tasks. Regression techniques, including linear regression and its variants, are widely used for predicting continuous outcomes based on the relationships between variables [James et al., 2013]. SVM is a powerful predictive method that aims to find the optimal hyperplane that best separates data points into different regions in a high-dimensional space [Cortes and Vapnik, 1995]. Bayesian classifiers are based on Bayes’ theorem and are probabilistic models that calculate the probability of a sample belonging to a particular class [Hand and Yu, 2001]. kNN is a simple yet effective algorithm that classifies a data point based on the majority class of its nearest neighbors in the feature space [Altman, 1992]. Elastic Net is a regularization technique combining L1 (Lasso) and L2 (Ridge) penalties, which is particularly useful for feature selection and dealing with multicollinearity in regression and classification problems [Zou and Hastie, 2005]. Each of these statistics-based methods has its own assumptions, strengths, and limitations, making them suitable for different types of datasets and prediction tasks.
In contrast, deep learning approaches like neural networks [Abiodun et al., 2018, Schmidhuber, 2014] and ResNet (Residual Network) [He et al., 2015] can learn complex relationships present in the data. Neural networks consist of layers of interconnected nodes that can model intricate patterns and interactions between variables, making them highly effective for both classification and prediction tasks. ResNet is a type of deep neural network that introduces skip connections to enable the training of very deep models by mitigating the vanishing gradient problem, allowing it to capture even more complex data relationships. These deep learning models are particularly powerful for handling large and high-dimensional datasets, offering a complementary approach to traditional statistics-based methods.
@@
In both prediction and classification tasks, various metrics are used to evaluate the performance of machine learning models. For the prediction tasks, we use common metrics like Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), Mean Squared Error (MSE), R-squared ($R^{2}$), Root Mean Squared Logarithmic Error (RMSLE), and Mean Absolute Percentage Error (MAPE). MAE measures the average magnitude of the errors between predicted and actual values. MSE is the average of the squared differences between predicted and actual values (RMSE is the square root of MSE). $R^{2}$ represents the proportion of the variance in the dependent variable that is predictable from the independent variables. RMSLE is particularly useful for tasks where the target variable spans several orders of magnitude. MAPE measures the percentage difference between predicted and actual values. For all the metrics, lower values indicate better fit whereas for $R^{2}$ values close to 1 are considered good.
@@
For classification tasks, we use metrics such as Accuracy, Precision, Recall, Area Under the Receiver Operating Characteristic Curve (AUC), F1 score, and Cohen’s Kappa. All these values can be derived from the confusion matrix. Accuracy measures the proportion of correctly classified instances over the total number of instances. Precision measures the proportion of true positive predictions among all positive predictions, focusing on the accuracy of positive predictions. Recall (also known as sensitivity) measures the proportion of true positive predictions among all actual positives, focusing on the model’s ability to detect positive instances. AUC measures the model’s ability to distinguish between the classes. F1 score is the harmonic mean of precision and recall, providing a balance between the two metrics. Cohen’s Kappa measures the agreement between predicted and actual classifications, accounting for the possibility of the agreement occurring by chance. For all the metrics, higher values indicate good performance.
@@
Feature analysis is performed using Feature Importance [Menze et al., 2009] and SHAP [Lundberg and Lee, 2017]. They provide valuable insights into how individual features impact the model and the outcomes. Feature Importance ranks the variables based on their contribution to the model’s performance, helping to identify which factors are most influential in the decision-making process. SHAP, on the other hand, explains how each feature impacts the outcome for each individual instance, highlighting both the global and the local effects. However, SHAP does not measure the contribution of a feature to the overall model performance but instead focuses on the direction and magnitude of a feature’s impact on specific outcomes. This dual approach enables a deeper understanding of the underlying dynamics in the data, allowing for more informed decisions and targeted improvements in complex systems like port operations.
In the context of predictive analytics for port operations, the aforementioned machine learning and deep learning methods are applied to predict and classify Total Time and Delay Time, which are critical factors in port logistics. By leveraging these techniques, we gain valuable insights into the intricate dynamics of port operations.
@@ 4 Results
We split the data (n=153331) into 80% (122664) for training and 20% (30667) for testing. We consider 24 (out of 35) features as independent variables and the other features related to calculation of Total Time and Delay Time at the port are removed. 17 of them are categorical variables. These categorical variables are transformed using one hot encoding (low cardinality variables like Month, Type of Mooring, Geographic Region, etc) or target encoding (high cardinality variables like Berth, Source, Destination, Port, etc). To ensure we don’t overfit and to pick the best hyperparameters, we perform 10-fold cross-validation (CV). We first look into the prediction of Total Time and Delay Time, followed by the classification of Total Time and Delay Time. The experiments were performed on an 8-core M2 CPU.
Also, we perform feature analysis for a comprehensive understanding of the key factors influencing port operations. Addressing these factors is crucial for optimizing the Berth Allocation Problem (BAP), Predictive Model Scheduling (PMS), port performance management, and risk assessment. By enhancing these aspects, ports can achieve greater sustainability, aligning with the goal of promoting green practices.
@@
Table 1 shows the performance of different Machine Learning and Deep Learning approaches for Total Time prediction. We see a general trend that tree-based methods are performing better than statistical approaches on this task. This might be because of the complex nature of the data and their superiority in handling categorical variables. Surprisingly, Neural network and ResNet performance is a bit underwhelming, the main reason for this is the presence of many categorical features in the data [Dahouda and Joe, 2021, Grinsztajn et al., 2022]. The High Cardinality, Non-Ordinal Nature, and Insufficient Sample Per Category are hampering the deep learning performance. We observe that Random Forest (RF) performs the best, has the lowest RMSE (10.4377), and also has better values in other metrics in comparison to the rest of the methods. Performance on the test set for RF is similar, where the RMSE is 10.2577 (other metrics are also in close range). This is visible from Fig. 2(a), where the predicted values and true values are near to the identity line with an $R^{2}$ value of 0.8287.
Fig. 2(b) shows the feature importance for the Total Time prediction, and we can observe that Berth feature plays the most crucial role due to some berths being more busy and less efficient compared to others, followed by other features like Moved Value, VL Containerized Load Weight, Type of Mooring Navigation, Source, Destination, Month, etc. Since, Berths play a vital role in port operations, efficient berth management is essential for optimizing port capacity, throughput, and overall smooth operations. The SHAP analysis can be seen in Fig. 3, where Fig. 3(a) indicates that features like Berth, Type of Mooring Navigation, and Nationality have a high impact on the outcomes for the random forest model. The horizontal spread indicates the magnitude of the effect that a feature has on the outcome whereas the color represents if the feature value is high or low. For example, the Berth feature (target encoded) denotes that, generally, a high value of this feature leads to a higher positive impact on the outcome. This means that it will lead to a higher Total Time value. Fig. 3(b) and Fig. 3(c) show the Decision and Force plot for a particular sample with true Total Time of 4.50 hours. The base value (49.52) represents the expected value of the model’s output before any features are considered. The model output is 4, which is near the truth. We see features like Berth, Type of Mooring Navigation, and Source have the highest impact on getting the right output for this sample. Since Berth and Source are fixed features, changing the Type of Mooring Navigation for this sample might help to reduce the Total Time. Note that for a different sample, some other features might be important.
(a) R-squared plot on test data for Total Time prediction
(b) Random Forest Feature importance for of Total Time prediction
Figure 2: Random Forest inference plots for Total Time prediction
(a) SHAP feature analysis plot for Total Time Prediction
(b) SHAP Decision plot for Total Time Prediction of a sample observation
(c) SHAP force plot for Total Time Prediction of a sample observation
Figure 3: SHAP Analysis for Total Time Prediction
We see a similar story from Table 2 where the tree-based methods are performing better for predicting the Delay Time at the port. Random Forest (RF) again performs the best with an RMSE of 4.2179. The test set performance for RF gives an RMSE of 4.3707 (other metrics are also in a similar range). From Fig. 4(a), we can see that the predicted and true values are not close to the identity line ($R^{2}$ is 0.6086). The prediction of Delay Time might be limited by the feature set, as it is known that weather information and other external factors play a key role with respect to delay.
Fig. 4(b) shows the feature importance for the Delay Time prediction, and we can observe that Moved Value, VL Weight Containerized Cargo, and Berth are key features followed by other features like Source, Destination, Month, etc. It is intuitive that vessels with higher Moved Value and VL Containerized Load Weight go through longer operations leading to build up in Delay Time. We can see the SHAP analysis in Fig. 5, where Fig. 5(a) shows that the Berth, Source, Month, and Moved Value have the most impact on the outcomes for the random forest model. Fig. 5(b) and Fig. 5(c) show the Decision and Force plot for a particular sample with true Delay Time of 16 hours. We get a predicted value of 15.17, which is very close to the truth, and can see that features like Month, Source, and Berth have the most impact towards getting the correct output for this sample. Unfortunately, for this sample, there is no potential for the outcome to be changed as the top features are fixed.
(a) R-squared plot on test data for Delay Time prediction
(b) Random Forest Feature importance for of Delay Time prediction
Figure 4: Random Forest inference plots for Delay Time prediction
(a) SHAP feature analysis plot for Delay Time Prediction
(b) SHAP Decision plot for Delay Time Prediction of a sample observation
(c) SHAP force plot for Delay Time Prediction of a sample observation
Figure 5: SHAP Analysis for Delay Time Prediction